# references.bib

### v1

@@ V Numerical Experiments
Example 1
This example shows that under Assumption 5, the $l_{1}$-norm estimator successfully recovers the true matrix, which supports Theorem 2. We generate a random matrix whose operator norm is 0.6. Whenever the attack times happen with some probability $p<1$, each coordinate $w_{t}^{i}$ of the disturbance $w_{t}$ at time $t$ is selected by the adversary to be $-\text{sgn}(x_{t}^{i})\cdot\gamma$, where $\gamma$ follows $\text{Uniform}[-3,-1]$ with probability 0.5 and $\text{Uniform}[10,20]$ with probability 0.5. The disturbances satisfy the symmetry assumption given in (20), but indeed have a nonzero mean.
@@
In Figure 1, we show the efficacy of the $l_{1}$-norm estimator compared to the least-squares and the $l_{2}$-norm estimator brought up in Section I. We report the error based on the Frobenius norm of the difference between the estimates and the true matrix. Note that all three estimators involve convex optimization, and therefore they can be efficiently solved by standard optimization solvers. We present various scenarios based on the attack probability $p$ and the dimension $d$. While the OLS and $l_{2}$-norm estimator behave in a similar pattern, with a plateau in the error and failure to identify the correct matrix, the $l_{1}$-norm estimator achieves the exact recovery for all scenarios. By comparing the figures with the same dimension, one can easily observe that as $p$ increases, indicating more frequent attacks, the time required for the exact recovery also increases. Similarly, by comparing the figures with the same $p$, it can be seen that a larger dimension leads to a slower exact recovery. These findings illustrate the recovery time of $O\bigr{(}\frac{\log 2d^{2}}{(1-p)^{2}}\bigr{)}$ stated in Theorem 2.
@@
Example 2
For this example, we generated a random matrix whose operator norm is 0.95. We compare the $l_{1}$-norm estimator with the other estimators in the case of completely arbitrary disturbances with the attack probability $p<0.5$, which supports Theorem 3. We choose each coordinate $w_{t}^{i}$ of the disturbance $w_{t}$ to be Gaussian with a mean of $100\cdot(\text{sgn}(x_{t}^{i})+2)$ and a variance of 5 when the system is under attack. This adversarial disturbance implies that whenever the attack times happen, the disturbance is almost always positive, so as to mislead the estimators towards positive bias. However, Figure 2 shows that the $l_{1}$-estimator successfully identify the matrix when $p$ is at most 0.47, while the other estimators again fails to improve beyond a certain level of error. Figures 2(a) and 2(b) imply that a slight change in $p$ leads to a considerable difference in the exact recovery time. This is because, unlike in Example 1, the recovery time is modified to $O\bigr{(}\frac{\log 2d^{2}}{(1-2p)^{2}}\bigr{)}$, which is very sensitive near $p=0.5$. Figure 2(c) demonstrates this sensitivity as $p$ approaches 0.5, showing that the error alternates between zero and positive values. Finally, when $p$ is exactly 0.5, Figure 2(d) indicates that the $l_{1}$-norm estimator has a plateau in error, similar to that of the other estimators. This aligns with Theorem 3, which only holds for $p<0.5$. This observation identifies the critical number 0.5 when the $l_{1}$-norm estimator ultimately recovers the true matrix even in the presence of arbitrarily large noncentral attacks.
@@ VI Conclusion
In this paper, we study the capability of the $l_{1}$-norm estimator to exactly identify the true matrix in the linear system identification problem, where the system suffers from adversarial noncentral disturbances. We show that the true matrix is exactly recovered under a symmetry assumption and the attack probability $p$ being less than 1. Furthermore, if $p<0.5$, the $l_{1}$-norm estimator prevails against any adversarial noncentral disturbances and achieves the exact recovery. This is the first result in the literature showing the possibility of accurate learning of systems under correlated, adversarial, and nonzero-mean disturbances.

### v5

@@ V Numerical Experiments
Example 1:
This example shows that under Assumption 5, the $l_{1}$-norm estimator successfully recovers the true matrix, which supports Theorem 2. We generate a random matrix whose operator norm is 0.6. Whenever the attack times happen with some probability $p<1$, the adversary selects each coordinate $w_{t}^{i}$ of the disturbance $w_{t}$ to be $-\text{sgn}(x_{t}^{i})\cdot\gamma$, where $\gamma$ follows $\text{Uniform}[-3,-1]$ with probability 0.5 and $\text{Uniform}[10,20]$ with probability 0.5. The disturbances satisfy the symmetry assumption given in (19), but indeed have a nonzero mean.
@@
In Figure 1, we show the efficacy of the $l_{1}$-norm estimator compared to the OLS and the $l_{2}$-norm estimator brought up in Section I. All three estimators involve convex optimization, and thus can be efficiently solved by standard optimization solvers. We report the error based on the Frobenius norm of the difference between the estimates and the true matrix. We present various scenarios based on the attack probability $p$ and the dimension $d$. While the OLS and $l_{2}$-norm estimator show a plateau in the error and fail to identify the correct matrix, the $l_{1}$-norm estimator achieves the exact recovery for all scenarios. The figures with the same dimension show that as $p$ increases towards 1, indicating more frequent attacks, the exact recovery time also increases. Similarly, the figures with the same $p$ demonstrate that a larger dimension leads to a slower exact recovery. These findings illustrate the recovery time of $\tilde{\Theta}\bigr{(}\frac{d}{p(1-p)^{2}}\bigr{)}$ stated in Theorem 2.
@@
Example 2:
For this example, we generated a random matrix whose operator norm is 0.95. We consider the case of completely arbitrary disturbances with the attack probability $p<0.5$, which supports Theorem 3. We choose each coordinate $w_{t}^{i}$ of the disturbance $w_{t}$ to be Gaussian with a mean of $100\cdot(\text{sgn}(x_{t}^{i})+2)$ and a variance of 5 when the system is under attack. This adversarial disturbance implies that the attack is almost always positive so as to mislead the estimators towards a positive bias. Figure 2 shows that the $l_{1}$-norm estimator successfully identifies the matrix when $p$ is at most 0.47, while the other estimators again fail to improve beyond a certain level of error. Figures 2(a) and 2(b) imply that a slight change in $p$ leads to a considerable difference in the exact recovery time. This is because, unlike in Example 1, the recovery time is modified to $\tilde{\Theta}\bigr{(}\frac{d}{p(1-2p)^{2}}\bigr{)}$, which is very sensitive near $p=0.5$. Figure 2(c) demonstrates this sensitivity as $p$ approaches 0.5, showing that the error alternates between zero and positive values. Finally, when $p$ is exactly 0.5, Figure 2(d) indicates that the $l_{1}$-norm estimator has a plateau in error, similar to that of the other estimators. This aligns with Theorem 3, which only holds for $p<0.5$. This observation identifies the critical bound of 0.5, below which the $l_{1}$-norm estimator can ultimately recover the true matrix even in the presence of arbitrarily large noncentral attacks.
@@ VI Conclusion
In this paper, we study the capability of the $l_{1}$-norm estimator to exactly identify the true matrix in the linear system identification problem, where the system suffers from adversarial noncentral disturbances. We show that the true matrix is exactly recovered under a symmetry assumption and the attack probability $p$ being less than 1. Furthermore, if $p<0.5$, the $l_{1}$-norm estimator prevails against arbitrary adversarial noncentral disturbances and achieves the exact recovery. This is the first result in the literature showing the possibility of accurate learning of systems under correlated, adversarial, and nonzero-mean disturbances.